# Gaudin models and moduli space of flower curves

### v1

@@ 2.3. Forests
$p:=\begin{cases}\text{the rightmost leaf on the left branch above $v$, }\quad\text{ if $v$ is internal}\\ \text{the rightmost leaf of $\tau$, }\quad\text{ if $v$ is the root}\end{cases}$
@@ 2.4. Families of subalgebras
Let $k$ be a commutative ring and let $W$ be a free finite-rank $k$ module. Let $X$ be a scheme over $k$. A family of $r$-dimensional subspaces of $W$ parametrized by $X$ is a rank $r$ locally free sheaf $\mathcal{V}\subset\mathcal{O}_{X}\otimes_{k}W$ such that $\mathcal{O}_{X}\otimes_{k}W/\mathcal{V}$ is locally free. Equivalently, we have a morphism of $k$-schemes $X\rightarrow{\mathrm{Gr}}(r,W)$. In particular, for each closed point $p\in X$, we have a dimension $r$ subspace
@@ 2.5. Algebras over $\mathbb{C}[\hbar]$
In this paper, we will often with algebras over $\mathbb{C}[\hbar]$. Given an algebra $R$ over $\mathbb{C}[\hbar]$, and $\varepsilon\in\mathbb{C}$, we let $R(\varepsilon):=R\otimes_{\mathbb{C}[\hbar]}\mathbb{C}[\hbar]/(\hbar-\varepsilon)$ be the specialization at $\hbar=\varepsilon$. Similarly, if $f:R\rightarrow S$ is a map of $\mathbb{C}[\hbar]$-algebras, we let $f(\varepsilon):R(\varepsilon)\rightarrow S(\varepsilon)$ be the map between their specializations. Similarly, for any scheme $X$ defined over $\mathbb{A}^{1}$, we write $X(\varepsilon)$ for the fibre over $\varepsilon\in\mathbb{C}$.
@@ 2.6. Quantum Hamiltonian reduction
The following special case will be particularly useful to us. Suppose that $\mathfrak{g}$ is a Lie algebra and that $\Phi:\mathfrak{g}\rightarrow A$ satisfies $\Phi([x,y])=[\Phi(x),\Phi(y)]$ for all $x,y\in\mathfrak{g}$ ($\Phi$ is called a quantum moment map). Suppose also that we are given a character $\theta:\mathfrak{g}\rightarrow\mathbb{C}$. Then we define $J\subset A$ to be the left ideal generated by $\{\Phi(x)-\theta(x):x\in\mathfrak{g}\}$.
@@
So it suffices to show that its image is exactly $(A/J)^{\mathfrak{g}}$. Let $[a]\in(A/J)^{\mathfrak{g}}$, let $b\in A$ and let $x\in\mathfrak{g}$. Then $\Phi(x)a-a\Phi(x)\in J$ which implies that $(\Phi(x)-\theta(x))a-a(\Phi(x)-\theta(x))\in J$. Multiplying by $b$ and using that $ba(\Phi(x)-\theta(x))\in J$, we conclude that $b(\Phi(x)-\theta(x))a\in J$. Since every element of $J$ is a sum of terms of the form $b(\Phi(x)-\theta(x))$, we conclude that $a\in N_{A}(J)$ as desired. ∎
We will be mostly concerned be concerned with the following situation. Let $\widetilde{\mathfrak{g}}$ be another Lie algebra with $\mathfrak{g}\subset\widetilde{\mathfrak{g}}$. Set $A=U\widetilde{\mathfrak{g}}$ and let $\Phi:\mathfrak{g}\rightarrow A$ be the obvious embedding. Suppose we are given a complementary subalgebra $\mathfrak{k}\subset\widetilde{\mathfrak{g}}$, so that $\widetilde{\mathfrak{g}}=\mathfrak{g}\oplus\mathfrak{k}$ as a vector space (but we do not assume that $\mathfrak{k}$ nor $\mathfrak{g}$ are ideals).
@@ 3.1. The Deligne-Mumford space and the cactus flower space
$M_{n+1}=\mathbb{C}^{n}\setminus\Delta/\mathbb{C}^{\times}\ltimes\mathbb{C}=(\mathbb{C}^{\times})^{n-1}\setminus\Delta/\mathbb{C}^{\times}$
@@
for distinct $i,j,k,l$.
There is a natural stratification of $\overline{M}_{n+1}$, indexed by $[n]$-labelled, rooted trees. Each stratum is a moduli space of nodal curves of the prescribed combinatorial type, which means we record the tree formed by the irreducible components and the labels of the marked points on each component. All these strata are locally closed subsets of $\overline{M}_{n+1}$ (they can be defined by imposing some $\mu_{ijk}=0$ and some $\mu_{ijk}\not=0$). The closure of each stratum is a union of strata of the same form. The dimension of a stratum is $n-k$, where $k$ is the number of internal vertices of the tree.
@@ 3.2. Moduli space of cactus flower curves
The moduli space $\overline{M}_{n+1}$ carries a line bundle $\widetilde{M}_{n+1}\rightarrow\overline{M}_{n+1}$ given by pulling back the tautological line bundle from ${\mathbb{P}}(\mathfrak{t}_{n})$, where $\mathfrak{t}_{n}=\mathbb{C}^{n}/\mathbb{C}$. A point of $\widetilde{M}_{n+1}$ is given by a framed cactus curve, which is a cactus curve $(C,{\underline{z}})$, along with a non-zero tangent vector $a\in T_{z_{n+1}}C$ at the “infinity” marked point, which we call the distinguished point. Note that this non-zero tangent vector has the effect of reducing the automorphism group of the component of $C$ containing $z_{n+1}$. In this way we get an embedding
$F_{n}:=\mathbb{C}^{n}\setminus\Delta/\mathbb{C}\rightarrow\widetilde{M}_{n+1}$
given by $(z_{1},\dots,z_{n})\mapsto({\mathbb{P}}^{1},(z_{1},\dots,z_{n},z_{n+1}=\infty),a=1)$. In our recent paper [IKLPR], we defined a compactification $\overline{F}_{n}$ of $F_{n}$ which contains $\widetilde{M}_{n+1}$ as a open subscheme. A point of $\overline{F}_{n}$ is a cactus flower curve, which is a curve $C=C_{1}\cup\dots\cup C_{m}$, where each $C_{j}$ is a framed cactus curve; these cactus curves are glued together at their distinguished point, and altogether carry $n$ distinct marked points.
More precisely, we construct $\overline{F}_{n}$ by starting with the matroid Schubert variety $\overline{\mathfrak{t}}_{n}$, [IKLPR, §3.2]. This is the subscheme of $\nu\in({\mathbb{P}}^{1})^{p([n])}$ defined by the equations
@@
for all distinct $i,j,k$. A point of $\overline{\mathfrak{t}}_{n}$ can be visualized as a flower curve, $C=C_{1}\cup\dots\cup C_{m}$, where each $C_{j}$ is a framed projective line, these are glued together at their distinguished point, and altogether carry $n$ (not necessarily distinct) marked points.
Equivalently, we can set $\delta_{ij}=\nu_{ij}^{-1}$. In these coordinates, the defining equations of $\overline{\mathfrak{t}}_{n}$ become
$\delta_{ij}+\delta_{jk}=\delta_{ik}\qquad\delta_{ij}+\delta_{ji}=0$
The matroid Schubert variety is stratified in two ways by set partitions ${\mathcal{S}}$ of $[n]$. Given set partitions ${\mathcal{S}},\mathcal{B}$, we have the strata
@@
There is a morphism $U_{\mathcal{S}}\rightarrow\prod_{k=1}^{m}{\mathcal{T}}_{S_{k}}$ and we define ${\widetilde{U}}_{\mathcal{S}}:=U_{\mathcal{S}}\times_{\prod_{k}\mathfrak{t}_{S_{k}}}\prod_{k}\widetilde{M}_{S_{k}+1}$.
@@ 3.3. Degeneration
Here the products of ${\mathbb{G}}$ are taken over $\mathbb{A}^{1}$; that is why there is only one $\varepsilon$ coordinate in ${\mathbb{G}}^{n}$. We let $\mathcal{F}_{n}={\mathbb{G}}^{n}\setminus\Delta/{\mathbb{G}}$ and ${\mathcal{T}}_{n}={\mathbb{G}}^{n}/{\mathbb{G}}$.
@@
We have isomorphisms $\mathcal{F}_{n}(\varepsilon)\cong M_{n+2}=(\mathbb{C}^{\times})^{n}\setminus\Delta/\mathbb{C}^{\times}$ for $\varepsilon\neq 0$ and $\mathcal{F}_{n}(0)\cong F_{n}=\mathbb{C}^{n}\setminus\Delta/\mathbb{C}$, defined by
@@
In [IKLPR], we defined fibrewise compactifications $\overline{\mathcal{T}}_{n}$ and ¯$\overline{\mathcal{F}}_{n}$ which fit into the diagram
@@
3.4. An open affine cover
We constructed ¯$\overline{\mathcal{F}}_{n}$ in [IKLPR] as follows. First, for each set partition ${\mathcal{S}}$, deforming $U_{S}$, we defined the open affine subscheme ${\mathcal{U}}_{\mathcal{S}}\subset\overline{\mathcal{T}}_{n}$,
Then we constructed $\widetilde{\mathcal{U}}_{\mathcal{S}}\subset(\mathbb{P}^{1})^{p([n])}\times\prod_{k}(\mathbb{P}^{1})^{t(S_{k})}\times\mathbb{C}$ a deformed version of ${\widetilde{U}}_{\mathcal{S}}$. Finally, we defined ¯$\overline{\mathcal{F}}_{n}$ by glueing these together. These schemes $\widetilde{\mathcal{U}}_{\mathcal{S}}$ are not in general affine.
${\mathcal{W}}_{\tau}=\{(\nu,\mu)\in\widetilde{\mathcal{U}}_{\mathcal{S}}:\mu_{ijk}\neq\infty\text{ if the meet of $i,k$ is above the meet of $i,j$ in $\tau$}\}$
@@ 3.4. An open affine cover
By a slight abuse of notation, we will write $\mu_{ikj}$ for the regular function on ${\mathcal{W}}_{\tau}$ which extends $\frac{z_{i}-z_{j}}{z_{i}-z_{k}}=\delta_{ji}\nu_{ki}$.
@@ 3.5. Commutative subalgebras of the Drinfeld-Kohno Lie algebra
(6) $[t_{ij},t_{kl}]=0\text{ for $i,j,k,l$ distinct}\quad[t_{ij},t_{ik}]=-[t_{ij},t_{jk}]\text{ for $i,j,k$ distinct}$
@@
These vectors satisfy $[h_{i}({\underline{z}}),h_{j}({\underline{z}})]=0$ and $\sum_{i}h_{i}({\underline{z}})=0$. Thus, they span an $n-1$-dimensional commutative subspace $Q({\underline{z}})\subset{\mathfrak{s}}_{n}^{1}$. It is easy to see that $Q({\underline{z}})$ only depends on the image of ${\underline{z}}$ in $M_{n+1}$. We let $N_{n}=\{Q({\underline{z}}):{\underline{z}}\in M_{n+1}\}\subset{\mathrm{Gr}}(n-1,{\mathfrak{s}}_{n}^{1})$ be the set of all subspaces constructed in this manner. Let $\overline{N}_{n}$ be the closure of this locus.
@@
Since commutativity is a closed condition, every $V\in\overline{N}_{n}$ is commutative. In fact, Aguirre-Felder-Veselov [AFV] showed that every maximal commutative subspace of ${\mathfrak{s}}_{n}^{1}$ has dimension $n-1$ and is contained in $\overline{N}_{n}$. Thus, we can regard $\overline{N}_{n}$ as the space of all maximal commutative subspaces of ${\mathfrak{s}}_{n}^{1}$.

### v2

@@ 2.3. Forests
[…]
@@ 2.4. Families of subalgebras
Let $k$ be a commutative ring and let $W$ be a free finite-rank $k$ module. Let $X$ be a scheme over $k$. A family of $r$-dimensional subspaces of $W$ parametrized by $X$ is a rank $r$ locally free sheaf $\mathcal{V}\subset\mathcal{O}_{X}\otimes_{k}W$ such that $(\mathcal{O}_{X}\otimes_{k}W)/\mathcal{V}$ is locally free. Equivalently, we have a morphism of $k$-schemes $X\rightarrow{\mathrm{Gr}}(r,W)$. In particular, for each closed point $p\in X$, we have a dimension $r$ subspace
@@ 2.5. Algebras over $\mathbb{C}[\hbar]$
In this paper, we will often work with algebras over $\mathbb{C}[\hbar]$. Given an algebra $R$ over $\mathbb{C}[\hbar]$, and $\varepsilon\in\mathbb{C}$, we let $R(\varepsilon):=R\otimes_{\mathbb{C}[\hbar]}\mathbb{C}[\hbar]/(\hbar-\varepsilon)$ be the specialization at $\hbar=\varepsilon$. Similarly, if $f:R\rightarrow S$ is a map of $\mathbb{C}[\hbar]$-algebras, we let $f(\varepsilon):R(\varepsilon)\rightarrow S(\varepsilon)$ be the map between their specializations. Similarly, for any scheme $X$ defined over $\mathbb{A}^{1}$, we write $X(\varepsilon)$ for the fibre over $\varepsilon\in\mathbb{C}$.
@@ 2.6. Quantum Hamiltonian reduction
The following special case will be particularly useful to us. Suppose that $\mathfrak{g}$ is a Lie algebra and that $\Phi:\mathfrak{g}\rightarrow A$ satisfies $\Phi([x,y])=[\Phi(x),\Phi(y)]$ for all $x,y\in\mathfrak{g}$ ($\Phi$ is called a quantum comoment map). Suppose also that we are given a character $\theta:\mathfrak{g}\rightarrow\mathbb{C}$. Then we define $J\subset A$ to be the left ideal generated by $\{\Phi(x)-\theta(x):x\in\mathfrak{g}\}$.
@@
So it suffices to show that its image is exactly $(A/J)^{\mathfrak{g}}$. Let $a]\in(A/J)^{\mathfrak{g}}$, let $b\in A$ and let $x\in\mathfrak{g}$. Then $\Phi(x)a-a\Phi(x)\in J$ which implies that $(\Phi(x)-\theta(x))a-a(\Phi(x)-\theta(x))\in J$. Multiplying by $b$ and using that $ba(\Phi(x)-\theta(x))\in J$, we conclude that $b(\Phi(x)-\theta(x))a\in J$. Since every element of $J$ is a sum of terms of the form $b(\Phi(x)-\theta(x))$, we conclude that $a\in N_{A}(J)$ as desired. ∎
We will be mostly concerned with the following situation. Let $\widetilde{\mathfrak{g}}$ be another Lie algebra with $\mathfrak{g}\subset\widetilde{\mathfrak{g}}$. Set $A=U\widetilde{\mathfrak{g}}$ and let $\Phi:\mathfrak{g}\rightarrow A$ be the obvious embedding. Suppose we are given a complementary subalgebra $\mathfrak{k}\subset\widetilde{\mathfrak{g}}$, so that $\widetilde{\mathfrak{g}}=\mathfrak{g}\oplus\mathfrak{k}$ as a vector space (but we do not assume that $\mathfrak{k}$ nor $\mathfrak{g}$ are ideals).
@@ 3.1. The Deligne-Mumford space and the cactus flower space
$M_{n+1}=(\mathbb{C}^{n}\setminus\Delta)/\mathbb{C}^{\times}\ltimes\mathbb{C}=((\mathbb{C}^{\times})^{n-1}\setminus\Delta)/\mathbb{C}^{\times}$
@@
for distinct $i,j,k,l$. (Here $t([n])$ is the set of triples, see §2.2.)
There is a natural stratification of $\overline{M}_{n+1}$, indexed by $[n]$-labelled, rooted trees. Each stratum is a moduli space of nodal curves of the prescribed combinatorial type, which means we record the tree formed by the irreducible components and the labels of the marked points on each component. All these strata are locally closed subsets of $\overline{M}_{n+1}$ (they can be defined by imposing some $\mu_{ijk}=0$ and some $\mu_{ijk}\not=0$). The closure of each stratum is a union of strata of the same form. The dimension of a stratum is nkn-k, where $k$ is the number of internal vertices of the tree.
@@ 3.2. Moduli space of cactus flower curves
The map $M_{n+1}=(\mathbb{C}^{n}\setminus\Delta)/\mathbb{C}^{\times}\ltimes\mathbb{C}\hookrightarrow{\mathbb{P}}(\mathfrak{t}_{n})$, where $\mathfrak{t}_{n}=\mathbb{C}^{n}/\mathbb{C}$, extends to a map $\overline{M}_{n+1}\rightarrow{\mathbb{P}}(\mathfrak{t}_{n})$ given by collapsing all components not containing the marked point $n+1$. Thus, the moduli space $\overline{M}_{n+1}$ carries a line bundle $\widetilde{M}_{n+1}\rightarrow\overline{M}_{n+1}$ given by pulling back the tautological line bundle from ${\mathbb{P}}(\mathfrak{t}_{n})$. A point of $\widetilde{M}_{n+1}$ is given by a framed cactus curve, which is a cactus curve $(C,{\underline{z}})$, along with a non-zero tangent vector $a\in T_{z_{n+1}}C$ at the “infinity” marked point, which we call the distinguished point (see [IKLPR, Remark 4.8]). Note that this non-zero tangent vector has the effect of reducing the automorphism group of the component of $C$ containing $z_{n+1}$. In this way we get an embedding
$F_{n}:=(\mathbb{C}^{n}\setminus\Delta)/\mathbb{C}\rightarrow\widetilde{M}_{n+1}$
given by $(z_{1},\dots,z_{n})\mapsto({\mathbb{P}}^{1},(z_{1},\dots,z_{n},z_{n+1}=\infty),a=1)$. In our recent paper [IKLPR], we defined a compactification $\overline{F}_{n}$ of $F_{n}$ which contains $\widetilde{M}_{n+1}$ as an open subscheme. A point of $\overline{F}_{n}$ is a cactus flower curve, which is a curve $C=C_{1}\cup\dots\cup C_{m}$, where each $C_{j}$ is a framed cactus curve; these cactus curves are glued together at their distinguished point, and altogether carry $n$ distinct marked points.
More precisely, we construct $\overline{F}_{n}$ by starting with the matroid Schubert variety $\overline{\mathfrak{t}}_{n}$, [IKLPR, §3.2]. On $\mathfrak{t}_{n}=\mathbb{C}^{n}/\mathbb{C}$, we have the natural functions $\delta_{ij}=z_{i}-z_{j}$, which collectively give a linear embedding $\mathfrak{t}_{n}\hookrightarrow\mathbb{C}^{p([n])}$, where $p([n])$ is the set of pairs, see §2.2. Then $\overline{\mathfrak{t}}_{n}$ is the closure of $\mathfrak{t}_{n}$ inside $({\mathbb{P}}^{1})^{p([n])}$.
More explicitly, $\overline{\mathfrak{t}}_{n}$ is the subscheme of $({\mathbb{P}}^{1})^{p([n])}$ defined by the equations
$\delta_{ij}+\delta_{jk}=\delta_{ik}\qquad\delta_{ij}+\delta_{ji}=0$
for distinct $i,j,k$. A point of $\overline{\mathfrak{t}}_{n}$ can be visualized as a flower curve, $C=C_{1}\cup\dots\cup C_{m}$, where each $C_{j}$ is a framed projective line, these are glued together at their distinguished point, and altogether carry $n$ (not necessarily distinct) marked points.
Equivalently, we can set $\nu_{ij}=\delta_{ij}^{-1}$. In these coordinates, the defining equations of $\overline{\mathfrak{t}}_{n}$ become
@@
The matroid Schubert variety is stratified in two ways by set partitions $\mathcal{S}$} of $[n]$. Given set partitions $\mathcal{S}$$\mathcal{B}$, we have the strata
@@
There is a morphism $U_{\mathcal{S}}\rightarrow\prod_{k=1}^{m}\mathfrak{t}_{S_{k}}$ and we define ${\widetilde{U}}_{\mathcal{S}}:=U_{\mathcal{S}}\times_{\prod_{k}\mathfrak{t}_{S_{k}}}\prod_{k}\widetilde{M}_{S_{k}+1}$. Here we use $\mathfrak{t}_{S}=\mathbb{C}^{S}/\mathbb{C}$ and $\widetilde{M}_{S+1}$ for the versions of these spaces where the marked points are labelled by the finite set $S$.
@@ 3.3. Degeneration
Here the products of $\mathbb{G}$} are taken over $\mathbb{A}^{1}$; that is why there is only one $\varepsilon$ coordinate in ${\mathbb{G}}^{n}$. We let $\mathcal{F}_{n}=({\mathbb{G}}^{n}\setminus\Delta)/{\mathbb{G}}$ and ${\mathcal{T}}_{n}={\mathbb{G}}^{n}/{\mathbb{G}}$.
@@
We have isomorphisms $\mathcal{F}_{n}(\varepsilon)\cong M_{n+2}=((\mathbb{C}^{\times})^{n}\setminus\Delta)/\mathbb{C}^{\times}$ for $\varepsilon\neq 0$ and $\mathcal{F}_{n}(0)\cong F_{n}=(\mathbb{C}^{n}\setminus\Delta)/\mathbb{C}$, defined by
@@
In [IKLPR], we defined fibrewise compactifications $\overline{{\mathcal{T}}}_{n}$ and $\overline{{\mathcal{F}}}_{n}$ which fit into the diagram
(5) $\overline{M}_{n+2}$$\overline{{\mathcal{F}}}_{n}$$\overline{F}_{n}$$\overline{T}_{n}$$\overline{{\mathcal{T}}}_{n}$$\overline{\mathfrak{t}}_{n}$$\{\varepsilon\neq 0\}$${\mathbb{A}}^{1}$$\{0\}$}
The variety $\overline{{\mathcal{T}}}_{n}$ is defined as the subscheme of $(\nu,\varepsilon)\in({\mathbb{P}}^{1})^{p([n])}\times{\mathbb{A}}^{1}$ defined by the equations
(6) $\varepsilon\nu_{ik}+\nu_{ij}\nu_{jk}=\nu_{ik}\nu_{jk}+\nu_{ij}\nu_{ik}\qquad\nu_{ij}+\nu_{ji}=\hbar$
for all distinct $i,j,k$.
3.4. An open affine cover
We constructed $\overline{{\mathcal{F}}}_{n}$ in [IKLPR] as follows. First, for each set partition $\mathcal{S}$}, deforming $U_{S}$, we defined the open affine subscheme ${\mathcal{U}}_{\mathcal{S}}\subset\overline{{\mathcal{T}}}_{n}$,
Then we constructed $\widetilde{\mathcal{U}}_{\mathcal{S}}\subset(\mathbb{P}^{1})^{p([n])}\times\prod_{k}(\mathbb{P}^{1})^{t(S_{k})}\times{\mathbb{A}}^{1}$ a deformed version of ${\widetilde{U}}_{\mathcal{S}}$. Finally, we defined $\overline{{\mathcal{F}}}_{n}$ by glueing these together. These schemes $\widetilde{\mathcal{U}}_{\mathcal{S}}$ are not in general affine.
@@ 3.4. An open affine cover
Figure 2. A planar binary forest $\tau$ and the corresponding cactus flower curve. On ${\mathcal{W}}_{\tau}$, we have the following regular functions: $\nu_{14},\nu_{24},\nu_{34},$ $\delta_{12},\delta_{13},\delta_{23},$ $\mu_{142},\mu_{143},\mu_{243},\mu_{132}$.
@@
By a slight abuse of notation, we will write $\mu_{ikj}$ for the regular function on ${\mathcal{W}}_{\tau}$ which extends $\frac{z_{i}-z_{j}}{z_{i}-z_{k}}=\delta_{ji}\nu_{ki}$. We note the relation $\mu_{ijk}\nu_{ki}=\nu_{ji}$. See Figure 2 for an example.
@@ 3.5. Commutative subalgebras of the Drinfeld-Kohno Lie algebra
(7) $t_{ij},t_{kl}]=0\text{ for $i,j,k,l$ distinct}\quad[t_{ij},t_{ik}]=-[t_{ij},t_{jk}]\text{ for $i,j,k$ distinct}$
@@
These vectors satisfy $h_{i}({\underline{z}}),h_{j}({\underline{z}})]=0$ and $\sum_{i}h_{i}({\underline{z}})=0$. Thus, they span an n1n-1-dimensional commutative subspace $Q({\underline{z}})\subset{\mathfrak{s}}_{n}^{1}$. It is easy to see that $Q({\underline{z}})$ only depends on the image of $\underline{z}$} in $M_{n+1}$. We let $N_{n}=\{Q({\underline{z}}):{\underline{z}}\in M_{n+1}\}\subset{\mathrm{Gr}}(n-1,{\mathfrak{s}}_{n}^{1})$ be the set of all subspaces constructed in this manner. Let $\overline{N}_{n}$ be the closure of this locus.
@@
Since commutativity is a closed condition, every $V\in\overline{N}_{n}$ is commutative. In fact, Aguirre-Felder-Veselov [AFV] showed that every maximal commutative subspace of ${\mathfrak{s}}_{n}^{1}$ has dimension n1n-1 and is contained in $\overline{N}_{n}$. Thus, we can regard $\overline{N}_{n}$ as the space of all maximal commutative subspaces of ${\mathfrak{s}}_{n}^{1}$.